Required information

Upriver Parts manufactures two products, V-1 and V-2, at its River Plant. Selected data for an average month for the two products follow.

V-1 V-2
Units produced 10,000 1,000
Direct materials cost per unit $ 2 $ 4
Machine hours per unit 1 2
Production runs per month 80 40


Production at the plant is automated and any labor cost is included in overhead. Data on manufacturing overhead at the plant follow.

Machine depreciation $ 57,000
Setup labor 26,400
Material handling 16,080
Total $ 99,480

Required:

a. Compute the unit costs for the two products V-1 and V-2 using the current costing system at Upriver (using machine hours as the allocation basis). (Do not round intermediate calculations. Round your answers to 2 decimal places.)
b. Compute the unit costs for the two products V-1 and V-2 using the proposed ABC system at Upriver. (Do not round intermediate calculations. Round your answers to 2 decimal places.)

Unit Cost
V-1 V-2
a. Using current costing system
b. Using proposed ABC system
Answer #1
Current System
V-1 V-2
Direct material 2 4
Overhead 8.29 16.58
Unit cost 10.29 20.58
ABC System
V-1 V-2
Direct material 2 4
Overheads
Machine Depreciation 47500 9500
Setup Labor 17600 8800
Material Handling 13400 2680
Total overheads 78500 20980
Overhead per unit 7.85 20.98
Unit cost 9.85 24.98

B c 1 Current System 2 V-1 V-2 4 =99480*1/12000 =SUM(B3:34) =99480*2/12000 =SUM(C3:04) V-1 12 V-2 4 3 Direct material 4 Overh